Calculate effective nuclear charge of Fe2+ ion by Slater's rule.

Grade:12th pass

1 Answers

REENU SIVADARSHINI.M
5 years ago
 EFFECTIVE NUCLEAR CHARGE = Z -S    Z= ATOMIC NO.
  S CAN BE CALCULATEDBY THIS METHOD 
S =[1s]        [ 2s 2p6]    [3s2 3p6 3d6 ]  [ fe+2=24 electrons]
      [n-2]           [n-1]                  [n]
 value of n=0.35;[n-1]=0.85 ;[n-2]=1;[n-3]=1
 now no. of electron * respective value
  s=[ 2*1] +[8*0.85]+[13*0.35]       {in last set reduce one electron  so 13 and not 14}
 s=2 + 6.80+4.55
s=13.35
 effective nuclear charge= 26- 13.35= 12.65